1.The pdf of a continuous random variable x is given by f(x)={c/x1/2 ,0<x<4 both inclusive

{0 ,elsewhere

find the value of the constant c,the cdf of x and p(x>or equal to 1)


1
Expert's answer
2021-07-13T05:06:15-0400
"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{4}\dfrac{c}{x^{1/2}}dx=c[2\sqrt{x}]\begin{matrix}
 4 \\
 0
\end{matrix}=4c=1"

"c=\dfrac{1}{4}"


"f(x)=\begin{cases}
 \dfrac{1}{4\sqrt{x}} & 0\leq x\leq 4\\
 0 & elsewhere
\end{cases}"


"F(x)=\displaystyle\int_{-\infin}^{x}f(x)dx, -\infin<x<\infin"


"F(x)=\begin{cases}
0, & x<0 \\
 \dfrac{\sqrt{x}}{2} & 0\leq x< 4\\
 1 & x\geq4
\end{cases}"

"P(x\geq1)=1-F(1)=1- \dfrac{\sqrt{1}}{2}= \dfrac{1}{2}"